# ComicsPAP: understanding comic strips by picking the correct panel

### v2

@@ 1 Introduction
Large multimodal models (LMMs) have achieved remarkable progress in understanding diverse data modalities, excelling at tasks such as image captioning [3, 1], visual question answering (VQA) [1, 2, 26, 3], and video comprehension [16]. However, these models still struggle with comprehending comic strips, which pose a unique challenge due to their distinct narrative structure. In video analysis, consecutive frames are visually similar and exhibit only minor changes over short intervals, which facilitates motion analysis. In contrast, comic panels are deliberately designed to depict discrete scenes of a story, resulting in significant visual differences from one panel to the next. Moreover, unlike videos—where segmentation into frames is inherent—comic pages require the reader (or an AI) to infer the segmentation, often demarcated by gutters that can vary in clarity and may not always be present. Additionally, textual and speech information is fragmented, embedded within speech bubbles or captions, and must be contextually integrated across non-sequential panels. Thus, effective comic understanding demands that models infer subtle inter-panel relationships, recognize recurring characters despite diverse artistic interpretations, and unravel causal sequences from sparse visual cues. These challenges render comic strips an ideal benchmark for advancing the next generation of vision-language models.
@@
In this work, we address the lack of comic strip analysis for LMMs by benchmarking foundational models for comic panel sequence understanding tasks. Our contributions are threefold: (i) we introduce the ComicsPAP dataset, structured around three distinct skills and five subtasks, which reframes sequence understanding as the task of selecting the appropriate panel that best fits within a given ordered set—an approach we term “Pick A Panel”; (ii) we evaluate state-of-the-art LMMs on ComicsPAP and demonstrate that their performance is no better than random choice; (iii) we fine-tune models with 3B and 7B parameters on ComicsPAP, achieving superior performance compared to larger models on this task.
The ComicsPAP dataset is publicly available on Hugging Face, as are the fine-tuned model weights. The training and evaluation code are available in the github repository, and an evaluation server is hosted on the competition page.
@@ 2 Related works
Availability and copyrighted data remain the foremost challenges in comics analysis. Many previous studies have addressed these issues by linking to original sources to collect images while only sharing the accompanying annotations [20, 18]. However, such approaches face significant limitations when scaled up. For instance, source material may become inaccessible over time, resulting in non-reproducible experimental settings (over 50% of LAION images [14] are no longer available). An alternative strategy involves providing scripts to download images directly from the original sources [13] and sharing only cropped versions [12]. Yet these methods also have drawbacks: sources can expire, and users might lack the necessary rights to download copyrighted images. To overcome these challenges, we exclusively utilize non-copyrighted, American-comic-styled materials from the renowned DCM website111https://digitalcomicmuseum.com/index.php. This choice enables us to share both the images and their annotations freely.
@@
Early work in panel sequence analysis, such as COMICS [7], introduced “closure” tasks where models select the correct subsequent panel from options based on visual cues, combined visual and textual features, or character-driven attributes. However, these methods relied on outdated OCR and panel extraction, resulting in poor-quality annotations and distractors sampled from unrelated pages—issues that rendered the task relatively trivial (with recent transformer approaches yielding already significant improvements). Subsequent work like ComicVT5 [21] improved OCR accuracy, sampled distractor panels from adjacent pages, and extended the task to text generation rather than classification. More recently, MangaUB [6] extended these evaluations to manga panels with tasks ranging from single-panel analysis (predicting weather, time, and counting characters) to comic strip understanding, yet even basic inferences proved challenging for LMMs. Complementing these efforts, ComiCap [19] showed that, with proper model selection, vision-language models can generate detailed captions in a zero-shot setting, underscoring the potential for further advances in comic panel comprehension.
Despite these advances, panel sequence analysis has largely stagnated with tasks based on outdated, automatically generated annotations for next-panel classification using randomly sampled alternatives. This approach suffers from several intrinsic limitations: (i) comics are seldom segmented into coherent stories before analysis, so randomly sampled panels often come from different narratives, styles, and character sets, making the task inherently easier; (ii) tasks are predominantly framed as next-panel prediction, neglecting the evaluation of causal reasoning through the reconstruction of sequences with missing panels; and (iii) the focus remains largely visual, with little emphasis on selecting the correct panel based solely on textual descriptions.
@@ 3 Dataset Creation
In this section, we describe the construction and annotation procedures of the ComicsPAP dataset. Our approach is driven by explicitly defining the cognitive skills we aim to assess and designing corresponding tasks that rigorously evaluate these skills. Specifically, we focus on three primary multimodal understanding skills—narrative anticipation, multi-frame reasoning, and co-reference resolution. Each skill is operationalized through carefully designed tasks under a unified evaluation paradigm we term “Pick A Panel”, where models select the correct panel given multimodal context and options.
@@
The narrative anticipation skill measures a model’s capability to infer immediate narrative continuation (or next-panel), similar to the next-token prediction task in NLP [11]. It relies on unidirectional contextual information derived from preceding panels, assessing the model’s understanding of narrative structure and visual-textual continuity. We adapt the well-established closure tasks from the COMICS dataset [7] into three refined variants: character coherence, visual closure, and text closure. All these tasks provide three consecutive context panels, requiring models to select the correct subsequent panel among distractors sampled from the same coherent narrative segment. This addresses a key limitation from prior work, where distractors came from different books, allowing potential stylistic or character-based shortcuts. In the visual closure variant, all context textboxes remain visible, and distractors (six total) are sampled within a challenging -panel window around the context panels, substantially increasing task difficulty compared to prior “hard” settings. The text closure variant further standardizes evaluation by dynamically rendering texts within bounding boxes. To maintain difficulty and fairness, preventing models from exploiting visual font-size cues, distractors in text closure tasks are pre-filtered by text length similarity with the correct one. Additionally, character coherence specifically selects sequences where the final panel contains exactly two prominent textboxes (spanning multiple panels only if at least of the textbox area is within one panel).
@@
Co-reference resolution involves correctly interpreting implicit references across comic panels. While prior work demonstrated success with explicit textual cues such as "the first page" or "the second image" [8], ComicsPAP emphasizes implicit co-reference by requiring models to link detailed captions with adjacent panels. This skill expects models to infer coherent relationships across characters, settings, or semantic continuity, especially during scene transitions [7]. To evaluate this skill, we introduce the caption relevance task, where models select the subsequent panel corresponding to a given caption from among the next five sequential panels in the narrative. Captions are generated by a specialized model (detailed in later sections) to ensure consistency. Following prior analysis [7], we recognize four major intra-panel transitions in comics 222We anticipate similar proportions as our dataset is randomly sampled from a subset (Comics100) of the original COMICS collection.: (i) action-to-action (), where the previous action continues in the next panel; (ii) subject-to-subject (), where the same subject appears with a closely related action; (iii) continued conversation (), where dialogue flows seamlessly between panels; and (iv) scene-to-scene (), where a complete change in setting occurs. The remaining panels, designated as moment-to-moment, represent the same exact moment, while less than are parallel transitions where both visual and textual content vary drastically. To mitigate straightforward linguistic cues—particularly in continued conversations—we explicitly exclude spoken text from generated captions.
@@ 3.1 Benchmark Annotations
Captions. For caption generation, we leveraged the state-of-the-art Molmo-72B model333https://huggingface.co/allenai/Molmo-72B-0924 [3], based on the Qwen2-72B language model444https://huggingface.co/Qwen/Qwen2-72B [25] and trained on the novel Pixmo dataset. The model was constrained to output up to 512 tokens per caption (with most captions being considerably shorter) and was prompted to exclude spoken or narration text.
@@ 5 Conclusions
We have presented ComicsPAP, a comprehensive benchmark designed to evaluate comic strip understanding in large multimodal models. By assembling a high-quality dataset through a combination of manual and automatic annotations, we have defined and assessed five key tasks—sequence filling, character coherence, visual closure, text closure, and caption relevance—that together probe the nuanced abilities required for comics comprehension. Our benchmarks thus serve a dual purpose: (i) to expose the inherent limitations of current LMMs in working with comic panel sequences, and (ii) to motivate the fine-tuning process using our automatically generated training set in order to improve these results. Our experimental results indicate that fine-tuning on our automatically annotated training set can boost model performance by more than over their zero-shot baselines, surpassing larger models.
@@
Despite promising results, several avenues remain open for future investigation. First, our fine-tuning strategy used uniform sampling across all skills, but adaptive sampling strategies accounting for task difficulty may improve performance. Second, prompts in our experiments were generic; designing task-specific prompts could help models better utilize contextual cues. Third, while our study focused on foundational LMMs, emerging specialized models—such as Magi [13], which provides graph-based representations of comic elements—could significantly enhance performance if integrated with Graph Neural Networks (GNNs). Finally, combining multiple smaller, task-specific models into an ensemble might yield superior results with fewer parameters than larger monolithic architectures. Exploring these directions promises to further advance multimodal comic comprehension.
@@ Acknowledgements
This paper has been supported by the Consolidated Research Group 2021 SGR 01559 from the Research and University Department of the Catalan Government and by project PID2023-146426NB-100 funded by MCIU/AEI/10.13039/ 501100011033 and FSE+. With the support of the FI SDUR predoctoral grant program from the Department of Research and Universities of the Generalitat de Catalunya and co-financing by the European Social Fund Plus (2024FISDU_00095).
We acknowledge EuroHPC Joint Undertaking for awarding project ID EHPC-AI-2024A02-057 access to MareNostrum5 at BSC, Spain.

### v3

@@ 1 Introduction
Large multimodal models (LMMs) have achieved remarkable progress in understanding diverse data modalities, excelling at tasks such as image captioning [4, 1], visual question answering (VQA) [1, 2, 27, 4], and video comprehension [17]. These success stories share a crucial advantage: the underlying data are naturally segmented (e.g., video frames) and exhibit strong temporal or spatial coherence, easing the burden of cross-frame reasoning. Comic strips, by contrast, deliberately disrupt these assumptions. Each panel depicts a distinct moment, producing abrupt visual shifts that break the smooth continuity exploited by video models. Furthermore, segmentation is implicit—readers infer panel boundaries from gutters whose presence, shape, and clarity are highly variable. Linguistic signals are likewise fragmented: dialogue and narration appear in localized bubbles or captions and must be stitched together across non-adjacent panels. Effective comic comprehension therefore hinges on three intertwined capabilities: (i) inferring latent panel boundaries, (ii) resolving recurring characters under diverse artistic styles, and (iii) reconstructing causal story lines from sparse, discontinuous multimodal evidence. These challenges position comics as a rigorous stress test for next-generation vision–language models.
@@
In this work, we address the lack of comic strip analysis by benchmarking foundational models for comic panel sequence understanding. Our contributions are threefold: (i) we introduce the ComicsPAP dataset, structured around three distinct skills and five subtasks, which reframes sequence understanding as the task of selecting the appropriate panel that best fits within a given ordered set—an approach we term “Pick A Panel”; (ii) we evaluate state-of-the-art LMMs on ComicsPAP and demonstrate that their performance is no better than random choice; (iii) we fine-tune models with 3B and 7B parameters on ComicsPAP, achieving superior performance compared to larger models on this task.
The ComicsPAP dataset is publicly available on HuggingFace111https://huggingface.co/datasets/VLR-CVC/ComicsPAP, as are the fine-tuned model weights222https://huggingface.co/VLR-CVC/Qwen2.5-VL-3B-Instruct-lora-ComicsPAP. The training and evaluation code are available in the GitHub repository333https://github.com/llabres/ComicsPAP, and an evaluation server is hosted on the competition page444https://rrc.cvc.uab.es/?ch=31&com=tasks.
@@ 2 Related works
Availability and copyrighted data remain the foremost challenges in comics analysis. Many previous studies have addressed these issues by linking to original sources to collect images while only sharing the accompanying annotations [21, 19]. However, such approaches face significant limitations when scaled up. For instance, source material may become inaccessible over time, resulting in non-reproducible experimental settings (over 50% of LAION images [15] are no longer available). An alternative strategy involves providing scripts to download images directly from the original sources [14] and sharing only cropped versions [13]. Yet these methods also have drawbacks: sources can expire, and users might lack the necessary rights to download copyrighted images. To overcome these challenges, we exclusively utilize non-copyrighted, American-comic-styled materials from the renowned DCM website555https://digitalcomicmuseum.com/index.php. This choice enables us to share both the images and their annotations freely.
@@
Early work in panel sequence analysis, such as COMICS [8], introduced “closure” tasks where models select the correct subsequent panel from options based on visual cues, combined visual and textual features, or character-driven attributes. However, these methods relied on old-generation OCR and panel extraction, resulting in poor-quality annotations and distractors sampled from unrelated pages—issues that rendered the task relatively trivial (with recent transformer approaches yielding already significant improvements). Subsequent work like ComicVT5 [22] improved OCR accuracy, sampled distractor panels from adjacent pages, and extended the task to text generation rather than classification. More recently, MangaUB [7] extended these evaluations to manga panels with tasks ranging from single-panel analysis (predicting weather, time, and counting characters) to comic strip understanding, yet even basic inferences proved challenging for LMMs. Complementing these efforts, ComiCap [20] showed that, with proper model selection, vision-language models can generate detailed captions in a zero-shot setting, underscoring the potential for further advances in comic panel comprehension.
Despite these advances, panel sequence analysis has largely stagnated with tasks based on earlier, automatically generated annotations for next-panel classification using randomly sampled alternatives. This approach suffers from several intrinsic limitations: (i) comics are seldom segmented into coherent stories before analysis, so randomly sampled panels often come from different narratives, styles, and character sets, making the task inherently easier; (ii) tasks are predominantly framed as next-panel prediction, neglecting the evaluation of causal reasoning through the reconstruction of sequences with missing panels; and (iii) the focus remains largely visual, with little emphasis on selecting the correct panel based solely on textual descriptions.
@@ 3 Dataset Creation
This section describes the construction and annotation procedures of the ComicsPAP dataset. Our approach is driven by explicitly defining the cognitive skills we aim to assess and designing corresponding tasks that rigorously evaluate these skills. Specifically, we focus on three primary multimodal understanding skills—narrative anticipation, multi-frame reasoning, and co-reference resolution.
3.1 Tasks definition
Each skill is operationalized through carefully designed tasks under a unified evaluation paradigm we term “Pick A Panel”, where models select the correct panel given multimodal context and options. Figure 2 provides a visualization of the five tasks.
@@ 3.1 Tasks definition
The narrative anticipation skill measures a model’s capability to infer immediate narrative continuation (or next-panel), similar to the next-token prediction task in NLP [12]. It relies on unidirectional contextual information derived from preceding panels, assessing the model’s understanding of narrative structure and visual-textual continuity. We adapt the well-established closure tasks from the COMICS dataset [8] into three refined variants: character coherence, visual closure, and text closure. In particular, visual closure predicts the next panel using only visual features (hiding text in candidate panels), text closure predicts the correct text in the next panel (showing the next panel with different options for the text), and character coherence predicts the text associated to each character (showing the next panel with varying options of character-text linking). All these tasks provide three consecutive context panels, requiring models to select the correct subsequent (4th) panel among distractors sampled from the same coherent narrative segment. This addresses a key limitation from prior work, where distractors came from different books, allowing potential stylistic or character-based shortcuts. In the visual closure variant, all context textboxes remain visible, and distractors (six total) are sampled within a challenging -panel window around the context panels, substantially increasing task difficulty compared to prior “hard” settings. The text closure variant further standardizes evaluation by dynamically rendering texts within bounding boxes. To maintain difficulty and fairness, preventing models from exploiting visual font-size cues, distractors in text closure tasks are pre-filtered by text length similarity with the correct one. Additionally, character coherence specifically selects sequences where the final panel contains exactly two prominent textboxes. A textbox is assigned to a panel if at least 80% of its area overlaps with the panel bounding box.
@@
Co-reference resolution involves correctly interpreting implicit references across comic panels, which means finding implicit references in consecutive panels. Following prior analysis [8, 3], we recognize four major intra-panel transitions in comics 666We anticipate similar proportions as our dataset is randomly sampled from a subset (Comics100) of the original COMICS collection.: (i) action-to-action (), where the previous action continues in the next panel; (ii) subject-to-subject (), where the same subject appears with a closely related action; (iii) continued conversation (), where dialogue flows seamlessly between panels; and (iv) scene-to-scene (), where a complete change in setting occurs. The remaining panels, designated as moment-to-moment, represent the same exact moment, while less than are parallel transitions where both visual and textual content vary drastically. Based on these intra-panel transitions, we want the model to find the correct panel given some sort of information about the previous one, not favoring any of the previously described transition types. Thus, we use text description: a detailed caption of the previous panel. To evaluate this skill, we introduce the caption relevance task, where models select the subsequent panel corresponding to a given caption from among the next five sequential panels in the narrative. Captions are generated by a specialized model (detailed in later sections) to ensure consistency. To mitigate straightforward linguistic cues—particularly in continued conversations—we explicitly exclude spoken text from generated captions. While prior work demonstrated success with explicit textual cues such as "the first page" or "the second image" [9], ComicsPAP emphasizes implicit co-reference by requiring models to link detailed captions with adjacent panels. This skill expects models to infer coherent relationships across characters, settings, or semantic continuity, especially during scene transitions [8].
@@ 3.2 Benchmark Annotations
Captions. For caption generation, we leveraged the state-of-the-art Molmo-72B model777https://huggingface.co/allenai/Molmo-72B-0924 [4], based on the Qwen2-72B language model888https://huggingface.co/Qwen/Qwen2-72B [26] and trained on the novel Pixmo dataset. The model was constrained to output up to 512 tokens per caption (with most captions being considerably shorter) and was prompted to exclude spoken or narration text.
@@ 5 Conclusions
We have presented ComicsPAP, a novel benchmark designed to evaluate comic strip understanding in large multimodal models. By assembling a high-quality dataset through a combination of manual and automatic annotations, we have defined and assessed five key tasks—sequence filling, character coherence, visual closure, text closure, and caption relevance—that together probe the nuanced abilities required for comics comprehension. Our benchmarks thus serve a dual purpose: (i) to expose the inherent limitations of current LMMs in working with comic panel sequences, and (ii) to motivate the fine-tuning process using our automatically generated training set in order to improve these results. Our experimental results indicate that fine-tuning on our automatically annotated training set can boost model performance by more than over their zero-shot baselines, surpassing larger models.
@@
Despite promising results, several avenues remain open for future investigation. First, our fine-tuning strategy used uniform sampling across all skills, but adaptive sampling strategies accounting for task difficulty may improve performance. Also, instead of a single model trained of multiple tasks, it could be interesting to see performances of multiple task-specific fine-tuned models. Second, prompts in our experiments were generic; designing task-specific prompts could help models better utilize contextual cues. Third, while our study focused on foundational LMMs, emerging specialized models—such as Magi [14], which provides graph-based representations of comic elements—could significantly enhance performance if integrated with Graph Neural Networks (GNNs). Fourth, combining multiple smaller, task-specific models into an ensemble might yield superior results with fewer parameters than larger monolithic architectures. Finally, we designed the tasks to leverage automatic annotations with some human refinement. However, we believe the quality of captions and OCR outputs can significantly influence model performance; thus, comparing earlier-generation extraction methods with newer or manually crafted alternatives remains an important direction for future work. Exploring these directions promises to further advance multimodal comic comprehension.
@@ Acknowledgements
This paper has been supported by the Consolidated Research Group 2021 SGR 01559 from the Research and University Department of the Catalan Government, and by project PID2023-146426NB-100 funded by MCIU/AEI/10.13039/ 501100011033 and FEDER, UE. With the support of the FI SDUR predoctoral grant program from the Department of Research and Universities of the Generalitat de Catalunya and co-financing by the European Social Fund Plus (2024FISDU_00095).
We acknowledge EuroHPC Joint Undertaking for awarding the project ID EHPC-AI-2024A02-057 access to MareNostrum5 at BSC, Spain.